### v1

# Structure-preserving approximation of the non-isothermal Cahn-Hilliard system

Abstract
We propose and analyze a structure-preserving approximation of the non-isothermal Cahn-Hilliard equation using conforming finite elements for the spatial discretization and a problem-specific mixed explicit-implicit approach for the temporal discretization. To ensure the preservation of structural properties, i.e. conservation of mass and internal energy as well as entropy production, we introduce a suitable variational formulation for the continuous problem, based on the entropy equation. Analytical findings are supported by numerical tests, including convergence analysis.
keywords:
finite elements , phase-field , structure-preserving approximation, non-isothermal , Cahn-Hilliard
journal: Applied Mathematics and Computation
[…]
[mainz]organization=Johannes Gutenberg-Universität,addressline=Staudingerweg 9, city=Mainz, postcode=55099, country=Germany
1 Introduction
The Cahn-Hilliard equation is well-known in physics to describe conservative phase transitions like magnetism [21] and spinodal decomposition [10, 5], particularly in binary systems like alloys or liquid mixtures. These phenomena typically involve the separation of two phases when the system is quenched below a critical temperature. In many applications, the system is forced into a critical regime, often under the assumption of constant temperature, i.e., an isothermal system. The impact of non-isothermal conditions in such processes have been considered in [27, 34, 35]. Cahn-Hilliard models have also been applied to tumor growth, as seen in isothermal models [17] and their non-isothermal counterparts [25]. Additionally, non-isothermal phase-field models are of interest in material science, particularly in non-isothermal sintering and powder bed fusion [41, 40, 31]. In these scenarios, the system can cross the critical temperature barrier locally in both directions, necessitating the study of non-isothermal variations of the Cahn-Hilliard equation. In this work we consider the non-isothermal Cahn-Hilliard equation (NCH) in the spirit of Pawlow and Alt [2] given by
$\partial_{t}\phi-\operatorname{div}(\mathbf{M}\nabla\tfrac{\mu}{\vartheta}-\mathbf{C}\nabla\tfrac{1}{\vartheta})$ $=0,$ (1)
$\tfrac{\mu}{\vartheta}+\gamma\Delta\phi-\partial_{\phi}\tfrac{f(\phi,\vartheta)}{\vartheta}$ $=0,$ (2)
$\partial_{t}e(\phi,\vartheta)-\operatorname{div}(\mathbf{C}\nabla\tfrac{\mu}{\vartheta}-\mathbf{K}\nabla\tfrac{1}{\vartheta})$ $=0.$ (3)
Here $\phi$ is the phase-field variable distinguishing between phase A, where $\phi=0$ and phase B, where $\phi=1$, $\tfrac{\mu}{\vartheta}$ is the chemical potential and $\vartheta$ is the temperature. The matrices $\mathbf{M},\mathbf{K},\mathbf{C}$ are the mobility matrix, heat conductivity matrix and a cross-coupling matrix. The above system is complemented with initial data and stated on a periodic space domain $\Omega$ and the finite time interval $(0,T)$. In addition the system is subjected to the second law of thermodynamics, i.e. entropy is non-decreasing in time. The connection of the chemical potential $\frac{\mu}{\vartheta}$, the internal energy density $e$ and the entropy density $s$ is given a Helmholtz free energy of the following form
$F(\phi,\vartheta):=\frac{\gamma\vartheta}{2}|\nabla\phi|^{2}+f(\phi,\vartheta).$ (4)
This allows us to show that the system satisfies the following structural properties dictated by thermodynamics.
$\int_{\Omega}\phi(t)=\int_{\Omega}\phi(0),\qquad\int_{\Omega}e(\phi(t),\vartheta(t))=\int_{\Omega}e(\phi(0),\vartheta(0)),$
$\int_{\Omega}s(\phi(t),\vartheta(t))=\int_{\Omega}s(\phi(0),\vartheta(0))+\int_{0}^{t}\mathcal{D}(\tfrac{\mu}{\vartheta},\tfrac{1}{\vartheta})$
with the entropy production
$\mathcal{D}(\tfrac{\mu}{\vartheta},\tfrac{1}{\vartheta})=\int_{\Omega}\begin{pmatrix}\nabla\frac{1}{\vartheta}&\nabla\frac{\mu}{\vartheta}\end{pmatrix}\begin{pmatrix}\mathbf{K}&-\mathbf{C}\\ -\mathbf{C}&\mathbf{M}\end{pmatrix}\begin{pmatrix}\nabla\frac{1}{\vartheta}\\ \nabla\frac{\mu}{\vartheta}\end{pmatrix}.$
Under usual compatibility assumptions of the Onsager matrix
$\mathbf{L}=\begin{pmatrix}\mathbf{K}&-\mathbf{C}\\ -\mathbf{C}&\mathbf{M}\end{pmatrix},$
which is assumed to be positive (semi)-definite, the total entropy is produced as time evolves.
In the literature there have been several different generalizations of phase-field models, non-conserved or conserved, for the non-isothermal setting; see [32, 24] for some general ansatz and a quite intensive overview. The well-known generalization for non-conserved phase-field models was introduced by Penrose and Five [33], while for a conserved phase-field model by Pawlow and Alt [2]. In the latter the system driving force is the entropy. Both models and related systems can also be derived within the well-known GENERIC or SNET formalism, see [18]. Modeling approaches based on the micro-force balance in the spirit of Gurtin [23] can be found in [30, 29] and the least action principle in [14]. Caginalp derived models with finite speed propagation of temperature, see [9, 13]. We may also note that these systems can be supplemented with (quasi)-incompressible flow, see [1, 15]. Colli et al. have studied several models from an analytical point of view, see [4, 12, 13]. Existence of weak solutions and further analysis for the model of Pawlow and Alt can be found in [3, 26]. Note that these results are the only available theoretical results dealing with the cross-coupling terms.
To numerically approximate such models, a purely implicit Euler method was proposed by [32], using the inverse temperature as the state variable. This approach was generalized to an energy-stable method by González et al. [20] using the average vector field methods [19]. Recently, Brunk et al. [8, 7] introduced structure-preserving approximations based on this transformation. The system under consideration possesses a gradient structure. For the Cahn-Hilliard and Allen-Cahn equations, common numerical methods include convex-concave splitting [38], the scalar auxiliary variable (SAV) approach [37, 28], and Energy Quadratisation (EQ) [11, 42]. The SAV and EQ methods introduce auxiliary variables that are either time-dependent or space-time-dependent, thereby relaxing the driving functional energy or entropy. In the case of the internal energy equation, the Energy Quadratisation is used in [22, 39].
The aim of this work is to study the variational properties of the NCH system (1)-(3) and its systematic discretisation in space and time using the standard temperature as state variable. The novelties of this manuscript are
  • Problem-adapted reformulation, choosing the entropy equation instead of the internal energy equation, to reveal the variational structure.
  • Structure-preserving approximation via suitable application of standard discretisations in space and time.
The manuscript is structured as follows. In Section 2, we introduce relevant notation and reformulate the system by exchanging the internal energy equation by its equivalent entropy equation. Afterwards, in Section 3, we will use the reformulated system to derive a variational formulation which is suitable for a direct approximation by a standard finite element method. In Sections 4 and 5, the fully discrete scheme using mixed explicit-implicit time discretisation and a convex-concave splitting of the driving potential is proposed and analysed. Finally, we present a convergence test and illustrate the evolution of the system through an application-inspired example in Section 6.
2 Preliminaries
Before we present our discretization method and main results in detail, let us briefly introduce our notation and main assumptions, and recall some basic facts.
2.1 Notation
The system (1)–(3) is investigated on a finite time interval $(0,T)$. To avoid the discussion of boundary conditions, we consider a spatially periodic setting, i.e.,
  1. (A1)
    $\Omega\subset\mathbb{R}^{d}$, $d=1,2,3$ is a cube and identified with the $d$-dimensional torus $\mathcal{T}^{d}$, i.e. periodic boundary conditions are applied. Moreover, functions on $\Omega$ are assumed to be periodic throughout the paper.
[…]
As usual, we denote by $L^{p}(a,b;X)$, $W^{k,p}(a,b;X)$, and $H^{k}(a,b;X)$, the Bochner spaces of integrable or differentiable functions on the time interval $(a,b)$ with values in some Banach space $X$. If $(a,b)=(0,T)$, we omit reference to the time interval and briefly write $L^{p}(X)$. The corresponding norms are denoted, e.g., by $\|\cdot\|_{L^{p}(X)}$ or $\|\cdot\|_{H^{k}(X)}$.
For later use we will introduce the following theorem from [16]:
Theorem 1.
Let N and M be two positive integers and V be defined as follows:
$V=\{(\mathbf{x},\mathbf{y})\in\mathbb{R}^{N}\times\mathbb{R}^{M}\leavevmode\nobreak\ \text{such that}\leavevmode\nobreak\ y>0\},$
where the notation $\mathbf{y}>c$ means that each component of $\mathbf{y}$ is greater than $c$, $c\in\mathbb{R}$. Let $\mathbf{b}\in\mathbb{R}^{N}\times\mathbb{R}^{M}$ and $\mathbf{g}$ and $\mathbf{G}$ be two continuous functions respectively from $V$ and $V\times[0,1]$ to $\mathbb{R}^{N}\times\mathbb{R}^{M}$ satisfying:
  1. (i)
    $\mathbf{G}(\cdot,1)=\mathbf{g}(\cdot);$
  2. (ii)
    $\forall\alpha\in[0,1]$, if $\mathbf{v}$ is such that $\mathbf{G}(\mathbf{v},\alpha)=\mathbf{b}$ then $\mathbf{v}\in W$ , where $W$ is defined as follows:
    $W=\{(\mathbf{x},\mathbf{y})\in\mathbb{R}^{N}\times\mathbb{R}^{M}\leavevmode\nobreak\ \text{s.t.}\leavevmode\nobreak\ \|\mathbf{x}\|<C_{1}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \varepsilon<\mathbf{y}<C_{2}\}$
[…]
  3. (iii)
    the topological degree of $\mathbf{G}(\cdot,0)$ with respect to $\mathbf{b}$ and $W$ is equal to $d_{0}\neq 0$.
Then the topological degree of $\mathbf{G}(\cdot,1)$ with respect to $\mathbf{b}$ and $W$ is also equal to $d_{0}\neq 0$. Consequently, there exists at least one solution $\mathbf{v}\in W$ such that $\mathbf{g}(\mathbf{v})=\mathbf{b}$.
2.2 Assumptions
In this subsection we collect the minimal assumptions for the rest of the manuscript.
  1. (A2)
    The interface parameter $\gamma\in\mathbb{R}$ is positive.
  2. (A3)
    The Onsager matrix $\mathbf{L}(\phi,\nabla\phi,\vartheta)=:\mathbf{L}\in\mathbb{R}^{2d\times 2d}$ is assumed to be positive definite. It means that for positive constants $\lambda_{0},\lambda_{1}>0$ we have
    $\lambda_{1}|\xi|^{2}\geq\mathbf{\xi}^{\top}\mathbf{L}\mathbf{\xi}\geq\lambda_{0}|\xi|^{2},\quad\forall\mathbf{\xi}\in\mathbb{R}^{2d}.$
  3. (A4)
    For a smooth driving potential $f(\cdot,\cdot):\mathbb{R}\times\mathbb{R}_{+}\to\mathbb{R}$ we assume that for every fixed $\phi$ the potential $f(\phi,\cdot):\mathbb{R}_{+}\to\mathbb{R}$ is concave and goes to infinity as $\vartheta\to 0$. Furthermore, for every fixed ϑ+italic-$\vartheta\in\mathbb{R}_{+}$ the potential $f(\cdot,\vartheta)$ can be decomposed in a convex and a concave part. This separation is denoted by $f_{\text{vex}},f_{\text{cav}}.$
2.3 Structural identities
On the formal level one can choose between several formulations of the temperature contribution which allows to switch between the internal energy $e$ and the entropy $s$.
Lemma 2.
Regular solutions $(\phi,\mu,\vartheta)$ to the non-isothermal Cahn-Hilliard system (1)-(3) satisfy
$\langle\partial_{t}\phi,1\rangle=\langle\partial_{t}e(\phi,\vartheta),1\rangle=0,\quad\langle\partial_{t}s(\phi,\vartheta),1\rangle=\mathcal{D}(\tfrac{\mu}{\vartheta},\tfrac{1}{\vartheta})\geq 0.$ (5)
Before considering the proof of Lemma 2 we first derive the entropy equation. Therefore we recall the identity $e=F+\vartheta s$ which after rearrangement yields $s=(e-F)/\vartheta$ and
$\partial_{t}s(\phi,\vartheta)$ $=\partial_{t}\tfrac{e(\phi,\vartheta)-F(\phi,\vartheta))}{\vartheta}=\tfrac{1}{\vartheta}\partial_{t}e(\phi,\vartheta)-\tfrac{1}{\vartheta}\partial_{t}F(\phi,\vartheta)-\tfrac{e(\phi,\vartheta)-F(\phi,\vartheta)}{\vartheta^{2}}\partial_{t}\vartheta$
$=\tfrac{1}{\vartheta}\partial_{t}e(\phi,\vartheta)-\tfrac{1}{\vartheta}\partial_{\phi}F(\phi,\vartheta)\partial_{t}\phi-\tfrac{1}{\vartheta}\partial_{\vartheta}F(\phi,\vartheta)\partial_{t}\vartheta-\tfrac{s(\phi,\vartheta)}{\vartheta}\partial_{t}\vartheta.$
Using the relation $s(\phi,\vartheta)=-\partial_{\vartheta}F(\phi,\vartheta)$ and substituting in (4) leads to
$\partial_{t}s(\phi,\vartheta)$ $=\tfrac{1}{\vartheta}\partial_{t}e(\phi,\vartheta)-\tfrac{1}{\vartheta}\partial_{\phi}f(\phi,\vartheta)\partial_{t}\phi-\gamma\nabla\phi\cdot\nabla\partial_{t}\phi.$
Insertion of the internal energy equation (3) and the definition of the chemical potential $\tfrac{\mu}{\vartheta}$, cf. (2) then yields
$\partial_{t}s(\phi,\vartheta)$ $=\tfrac{1}{\vartheta}\operatorname{div}(\mathbf{C}\nabla\tfrac{\mu}{\vartheta}-\mathbf{K}\nabla\tfrac{1}{\vartheta})-\tfrac{\mu}{\vartheta}\partial_{t}\phi-\gamma\operatorname{div}(\partial_{t}\phi\nabla\phi).$
Now inserting the Cahn-Hilliard equation cf. (1) implies
$\partial_{t}s(\phi,\vartheta)$ $=\tfrac{1}{\vartheta}\operatorname{div}(\mathbf{C}\nabla\tfrac{\mu}{\vartheta}-\mathbf{K}\nabla\tfrac{1}{\vartheta})-\tfrac{\mu}{\vartheta}\operatorname{div}(\mathbf{M}\nabla\tfrac{\mu}{\vartheta}-\mathbf{C}\nabla\tfrac{1}{\vartheta})$
$\quad-\gamma\operatorname{div}(\partial_{t}\phi\nabla\phi).$
By using the reverse product rule we finally obtain the entropy equation
$\partial_{t}s(\phi,\vartheta)$ $=\nabla\tfrac{1}{\vartheta}\cdot\mathbf{K}\nabla\tfrac{1}{\vartheta}-2\nabla\tfrac{1}{\vartheta}\cdot\mathbf{C}\nabla\tfrac{\mu}{\vartheta}+\nabla\tfrac{\mu}{\vartheta}\cdot\mathbf{M}\nabla\tfrac{\mu}{\vartheta}$
$\quad-\operatorname{div}\left(\gamma\partial_{t}\phi\nabla\phi+\tfrac{1}{\vartheta}\mathbf{K}\nabla\tfrac{1}{\vartheta}-\tfrac{1}{\vartheta}\mathbf{C}\nabla\tfrac{\mu}{\vartheta}-\tfrac{\mu}{\vartheta}\mathbf{C}\nabla\tfrac{1}{\vartheta}+\tfrac{\mu}{\vartheta}\mathbf{M}\nabla\tfrac{\mu}{\vartheta}\right).$ (6)
Proof of Lemma 2.
Conservation of mass and internal energy as well as entropy production simply follow from the respective equations (1),(3) and (2.3) by integration over the domain and partial integration. ∎
Note that to obtain the entropy production for system (1)–(3) we need to test with $\tfrac{1}{\vartheta}$ which is not directly amendable for discretizations via finite elements. However, using the entropy equation instead of the internal energy equation implies that we only need to test with $\vartheta$ to get the energy conservation, while entropy production follows by simple integration over the domain and using partial integration for the flux.
System (1),(2),(2.3) is equivalent to (1)–(3) and therefore also satisfies conservation of mass and internal energy as well as entropy production.
3 Variational structure
In the this section we will reformulate system (1), (2), (2.3) into a variational formulation suitably such that conservation of mass and energy as well as entropy production follows directly by insertion of suitable test functions.
Lemma 3.
Regular solutions $(\phi,\mu,\vartheta)$ to system (1), (2), (2.3) satisfy the variational formulation
$\langle\partial_{t}\phi,\psi\rangle$ $=\langle\mathbf{M}\tfrac{\mu}{\vartheta^{2}}\nabla\vartheta-\tfrac{\mathbf{M}}{\vartheta}\nabla\mu-\tfrac{\mathbf{C}}{\vartheta^{2}}\nabla\vartheta,\nabla\psi\rangle$ (7)
$\langle\mu,\xi\rangle$ $=\gamma\langle\nabla\phi,\vartheta\nabla\xi+\xi\nabla\vartheta\rangle+\langle\partial_{\phi}f,\xi\rangle$ (8)
$\langle\partial_{t}s,\omega\rangle$ $=\langle\tfrac{\mathbf{K}}{\vartheta^{2}}\nabla\vartheta+\tfrac{\mathbf{C}}{\vartheta}\nabla\mu-2\mathbf{C}\tfrac{\mu}{\vartheta^{2}}\nabla\vartheta-\mathbf{M}\tfrac{\mu}{\vartheta}\nabla\mu+\mathbf{M}\left(\tfrac{\mu}{\vartheta}\right)^{2}\nabla\vartheta,\tfrac{\omega}{\vartheta^{2}}\nabla\vartheta-\tfrac{1}{\vartheta}\nabla\omega\rangle$
$\quad+\langle\tfrac{\mathbf{C}}{\vartheta^{2}}\nabla\vartheta+\tfrac{\mathbf{M}}{\vartheta}\nabla\mu-\mathbf{M}\tfrac{\mu}{\vartheta^{2}}\nabla\vartheta,\tfrac{\omega}{\vartheta}\nabla\mu\rangle+\gamma\langle\partial_{t}\phi\nabla\phi,\nabla\omega\rangle$ (9)
for smooth test functions $\psi,\xi,\omega.$
Proof.
First we multiply the equations (1), (2) and (2.3) by the functions $\psi,\vartheta\xi$ and $\omega$, respectively, and integrate over the domain $\Omega$ to obtain
$\langle\partial_{t}\phi,\psi\rangle$ $=\langle\operatorname{div}(\mathbf{M}\nabla\tfrac{\mu}{\vartheta}-\mathbf{C}\nabla\tfrac{1}{\vartheta}),\psi\rangle$
$\langle\mu,\xi\rangle$ $=-\gamma\langle\Delta\phi,\vartheta\xi\rangle+\langle\partial_{\phi}f,\xi\rangle$
$\langle\partial_{t}s,\omega\rangle$ $=\langle\nabla\tfrac{1}{\vartheta}\cdot\mathbf{K}\nabla\tfrac{1}{\vartheta}-2\nabla\tfrac{1}{\vartheta}\cdot\mathbf{C}\nabla\tfrac{\mu}{\vartheta}+\nabla\tfrac{\mu}{\vartheta}\cdot\mathbf{M}\nabla\tfrac{\mu}{\vartheta},\omega\rangle$
$\quad-\langle\operatorname{div}\left(\gamma\partial_{t}\phi\nabla\phi+\tfrac{1}{\vartheta}\mathbf{K}\nabla\tfrac{1}{\vartheta}-\tfrac{1}{\vartheta}\mathbf{C}\nabla\tfrac{\mu}{\vartheta}-\tfrac{\mu}{\vartheta}\mathbf{C}\nabla\tfrac{1}{\vartheta}+\tfrac{\mu}{\vartheta}\mathbf{M}\nabla\tfrac{\mu}{\vartheta}\right),\omega\rangle.$
In the first equation we use partial integration and derive
$\langle\partial_{t}\phi,\psi\rangle$ $=-\langle\mathbf{M}\nabla\tfrac{\mu}{\vartheta}-\mathbf{C}\nabla\tfrac{1}{\vartheta},\nabla\psi\rangle.$
Applying the chain rule to $\nabla\tfrac{\mu}{\vartheta}$ and $\nabla\tfrac{1}{\vartheta}$ we immediately get (7). In the second equation we use partial integration
$\langle\mu,\xi\rangle$ $=\gamma\langle\nabla\phi,\nabla(\xi\vartheta)\rangle+\langle\partial_{\phi}f,\xi\rangle$
and apply the product rule to get (8). Finally, in the third equation we again use partial integration
$\langle\partial_{t}s,\omega\rangle$ $=\langle\nabla\tfrac{1}{\vartheta}\cdot\mathbf{K}\nabla\tfrac{1}{\vartheta}-2\nabla\tfrac{1}{\vartheta}\cdot\mathbf{C}\nabla\tfrac{\mu}{\vartheta}+\nabla\tfrac{\mu}{\vartheta}\cdot\mathbf{M}\nabla\tfrac{\mu}{\vartheta},\omega\rangle$
$\quad+\langle\gamma\partial_{t}\phi\nabla\phi+\tfrac{1}{\vartheta}\mathbf{K}\nabla\tfrac{1}{\vartheta}-\tfrac{1}{\vartheta}\mathbf{C}\nabla\tfrac{\mu}{\vartheta}-\tfrac{\mu}{\vartheta}\mathbf{C}\nabla\tfrac{1}{\vartheta}+\tfrac{\mu}{\vartheta}\mathbf{M}\nabla\tfrac{\mu}{\vartheta},\nabla\omega\rangle.$
Combining with the chain rule for the fraction terms we get
$\langle\partial_{t}s,\omega\rangle$ $=\langle\tfrac{\mathbf{K}}{\vartheta^{2}}\nabla\vartheta+2\tfrac{\mathbf{C}}{\vartheta}\nabla\mu-2\mathbf{C}\tfrac{\mu}{\vartheta^{2}}\nabla\vartheta-2\mathbf{M}\tfrac{\mu}{\vartheta}\nabla\mu+\mathbf{M}\left(\tfrac{\mu}{\vartheta}\right)^{2}\nabla\vartheta,\tfrac{\omega}{\vartheta^{2}}\nabla\vartheta\rangle$
$\quad+\langle\tfrac{\mathbf{M}}{\vartheta}\nabla\mu,\tfrac{\omega}{\vartheta}\nabla\mu\rangle+\langle\gamma\partial_{t}\phi\nabla\phi,\nabla\omega\rangle$
$\quad+\langle-\tfrac{\mathbf{K}}{\vartheta^{2}}\nabla\vartheta-\tfrac{\mathbf{C}}{\vartheta}\nabla\mu+2\mathbf{C}\tfrac{\mu}{\vartheta^{2}}\nabla\vartheta+\mathbf{M}\tfrac{\mu}{\vartheta}\nabla\mu-\mathbf{M}\left(\tfrac{\mu}{\vartheta}\right)^{2}\nabla\vartheta,\tfrac{1}{\vartheta}\nabla\omega\rangle,$
which after rearrangement yields (9). ∎
Theorem 4.
Regular solutions $(\phi,\mu,\vartheta)$ to the variational formulation (7)–(9) satisfy
$\langle\partial_{t}\phi,1\rangle=\langle\partial_{t}e,1\rangle=0,\quad\langle\partial_{t}s,1\rangle=\mathcal{D}(\tfrac{\mu}{\vartheta},\tfrac{1}{\vartheta})\geq 0.$
Moreover, these properties follow directly by using test functions $\psi\in\{1,\mu\},\xi=\partial_{t}\phi,w\in\{1,\vartheta\}.$
Proof.
By setting $\psi=1$ we immediately find
$\langle\partial_{t}\phi,1\rangle=\langle\mathbf{M}\tfrac{\mu}{\vartheta^{2}}\nabla\vartheta-\tfrac{\mathbf{M}}{\vartheta}\nabla\mu-\tfrac{\mathbf{C}}{\vartheta^{2}}\nabla\vartheta,\nabla 1\rangle=0.$
The entropy production follows by using the test function $\omega=1$:
$\langle\partial_{t}s,1\rangle$ $=\langle\tfrac{\mathbf{K}}{\vartheta^{2}}\nabla\vartheta+\tfrac{\mathbf{C}}{\vartheta}\nabla\mu-2\mathbf{C}\tfrac{\mu}{\vartheta^{2}}\nabla\vartheta-\mathbf{M}\tfrac{\mu}{\vartheta}\nabla\mu+\mathbf{M}\left(\tfrac{\mu}{\vartheta}\right)^{2}\nabla\vartheta,\tfrac{1}{\vartheta^{2}}\nabla\vartheta\rangle$
$\quad+\langle\tfrac{\mathbf{C}}{\vartheta^{2}}\nabla\vartheta+\tfrac{\mathbf{M}}{\vartheta}\nabla\mu-\mathbf{M}\tfrac{\mu}{\vartheta^{2}}\nabla\vartheta,\tfrac{1}{\vartheta}\nabla\mu\rangle$
$=\langle\mathbf{K}\tfrac{1}{\vartheta^{2}}\nabla\vartheta,\tfrac{1}{\vartheta^{2}}\nabla\vartheta\rangle-2\langle\mathbf{C}\left(\tfrac{\mu}{\vartheta^{2}}\nabla\vartheta-\tfrac{1}{\vartheta}\nabla\mu\right),\tfrac{1}{\vartheta^{2}}\nabla\vartheta\rangle$
$\quad+\langle\mathbf{M}\left(\tfrac{\mu}{\vartheta^{2}}\nabla\vartheta-\tfrac{1}{\vartheta}\nabla\mu\right),\tfrac{\mu}{\vartheta^{2}}\nabla\vartheta-\tfrac{1}{\vartheta}\nabla\mu\rangle$
$=\langle\mathbf{K}\nabla\tfrac{1}{\vartheta},\nabla\tfrac{1}{\vartheta}\rangle-2\langle\mathbf{C}\nabla\tfrac{\mu}{\vartheta},\nabla\tfrac{1}{\vartheta}\rangle+\langle\mathbf{M}\nabla\tfrac{\mu}{\vartheta},\nabla\tfrac{\mu}{\vartheta}\rangle$
$=\mathcal{D}(\tfrac{\mu}{\vartheta},\tfrac{1}{\vartheta})\geq 0.$
Using the identities $e=F+\vartheta s$ and $s=-\partial_{\vartheta}F$, energy conservation follows:
$\langle\partial_{t}e,1\rangle$ $=\langle\partial_{t}(F+\vartheta s),1\rangle$
$=\langle\partial_{\phi}F\partial_{t}\phi+\partial_{\vartheta}F\partial_{t}\vartheta+\partial_{t}\vartheta s+\vartheta\partial_{t}s,1\rangle$
$=\langle\gamma\vartheta\nabla\phi\cdot\nabla\partial_{t}\phi+\partial_{\phi}f\partial_{t}\phi+\vartheta\partial_{t}s,1\rangle$
$=\gamma\langle\nabla\phi,\vartheta\nabla\partial_{t}\phi\rangle+\langle\partial_{\phi}f,\partial_{t}\phi\rangle+\langle\partial_{t}s,\vartheta\rangle.$
Taking $\xi=\partial_{t}\phi$ in (8) and $\omega=\vartheta$ in (9) we further get
$\langle\partial_{t}e,1\rangle$ $=\langle\mu,\partial_{t}\phi\rangle-\gamma\langle\nabla\phi,\partial_{t}\phi\nabla\vartheta\rangle+\gamma\langle\partial_{t}\phi\nabla\phi,\nabla\vartheta\rangle$
$\quad+\langle\tfrac{\mathbf{C}}{\vartheta^{2}}\nabla\vartheta+\tfrac{\mathbf{M}}{\vartheta}\nabla\mu-\mathbf{M}\tfrac{\mu}{\vartheta^{2}}\nabla\vartheta,\nabla\mu\rangle$
and using (7) with $\psi=\mu$ we finally obtain
$\langle\partial_{t}e,1\rangle$ $=\langle\mathbf{M}\tfrac{\mu}{\vartheta^{2}}\nabla\vartheta-\tfrac{\mathbf{M}}{\vartheta}\nabla\mu-\tfrac{\mathbf{C}}{\vartheta^{2}}\nabla\vartheta,\nabla\mu\rangle$
$\quad+\langle\tfrac{\mathbf{C}}{\vartheta^{2}}\nabla\vartheta+\tfrac{\mathbf{M}}{\vartheta}\nabla\mu-\mathbf{M}\tfrac{\mu}{\vartheta^{2}}\nabla\vartheta,\nabla\mu\rangle$
$=0.$
4 Fully discrete scheme for the non-isothermal Cahn-Hilliard equation
As preparatory step, let us introduce the relevant notation and assumptions for our discretization strategy.
Time discretization
We partition the time interval $[0,T]$ into uniformly divided subintervals with step size $\tau>0$ and introduce $\mathcal{I}_{\tau}:=\{0=t^{0},t^{1}=\tau,\ldots,t^{n_{\tau}}=T\}$, where $n_{\tau}=\tfrac{T}{\tau}$ is the absolute number of time steps. We denote by $\Pi^{1}_{c}(\mathcal{I}_{\tau}),\Pi^{0}(\mathcal{I}_{\tau})$ the spaces of continuous piecewise linear and piecewise constant functions on $\mathcal{I}_{\tau}$. By $g^{n+1},g^{n},g^{*}$ we denote the evaluation of $g\in\Pi^{1}_{c}(\mathcal{I}_{\tau}),\Pi^{0}(\mathcal{I}_{\tau})$ at $t=\{t^{n+1},t^{n},t^{*}\}$, where $t^{*}$ serves as a placeholder for any fixed $t^{*}$ with $t^{*}\in[t^{n},t^{n+1}]$. Further, we introduce the time difference and the discrete time derivative via
$d^{n+1}g=g^{n+1}-g^{n},\qquad d^{n+1}_{\tau}g=\frac{g^{n+1}-g^{n}}{\tau},$
respectively.
Space discretization
For the spacial discretisation we require that the mesh $\mathcal{T}_{h}$ is a geometrically conforming partition of $\Omega$ into simplices of maximal diameter $h$ that can be extended periodically to periodic extensions of $\Omega$. We denote the space of continuous and linear finite element functions over the mesh $\mathcal{T}_{h}$ via
$\mathcal{V}_{h}$ :=$:=\{v\in H^{1}(\Omega)\cap C^{0}(\bar{\Omega}):v|_{K}\in P_{1}(K),\leavevmode\nobreak\ \forall K\in\mathcal{T}_{h}\}.$
We denote the convex-concave splitting of the potential $f$ by
$f_{\phi}(\phi_{h}^{n+1},\phi_{h}^{n},\vartheta_{h}^{n+1}):=\partial_{\phi}f_{\text{vex}}(\phi_{h}^{n+1},\vartheta_{h}^{n+1})+\partial_{\phi}f_{\text{cav}}(\phi_{h}^{n},\vartheta_{h}^{n+1}).$
Fully discrete scheme
We propose a fully discrete time-stepping method for the non-isothermal Cahn-Hilliard system (1),(2),(2.3).
Problem 5.
Let the initial data $(\phi_{h}^{0},\vartheta_{h}^{0})\in\mathcal{V}_{h}\times\mathcal{V}_{h}$ be given. Find $(\phi_{h},\vartheta_{h})\in\Pi^{1}_{c}(\mathcal{I}_{\tau};\mathcal{V}_{h}\times\mathcal{V}_{h})$ and $\mu_{h}\in\Pi^{0}(\mathcal{I}_{\tau};\mathcal{V}_{h})$ such that
$\langle d^{n+1}_{\tau}\phi_{h},\psi_{h}\rangle$ $=\langle\mathbf{M}_{h}^{*}\tfrac{\mu_{h}^{*}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1}-\tfrac{\mathbf{M}_{h}^{*}}{\vartheta_{h}^{n+1}}\nabla\mu_{h}^{n+1}-\tfrac{\mathbf{C}_{h}^{*}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1},\nabla\psi_{h}\rangle$ (10)
$\langle\mu_{h}^{n+1},\xi_{h}\rangle$ $=\gamma\langle\nabla\phi_{h}^{n+1},\vartheta_{h}^{n+1}\nabla\xi_{h}\rangle+\gamma\langle\nabla\phi_{h}^{*},\xi_{h}\nabla\vartheta_{h}^{n+1}\rangle$
$\quad+\langle f_{\phi}(\phi_{h}^{n+1},\phi_{h}^{n},\vartheta_{h}^{n+1}),\xi_{h}\rangle$ (11)
$\langle d^{n+1}_{\tau}s_{h},\omega_{h}\rangle$ $=\langle\tfrac{K_{h}^{*}}{\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1}+\mathbf{C}_{h}^{*}\nabla\mu_{h}^{n+1}-2\mathbf{C}_{h}^{*}\tfrac{\mu_{h}^{*}}{\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1}-\mathbf{M}_{h}^{*}\mu_{h}^{*}\nabla\mu_{h}^{n+1}$
$\quad+\mathbf{M}_{h}^{*}\tfrac{\mu_{h}^{*}\mu_{h}^{*}}{\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1},\tfrac{\omega_{h}}{\vartheta_{h}^{n+1}\vartheta_{h}^{n+1}\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1}-\tfrac{1}{\vartheta_{h}^{*}\vartheta_{h}^{n+1}}\nabla\omega_{h}\rangle$
$\quad+\langle\tfrac{\mathbf{C}_{h}^{*}}{\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1}+\mathbf{M}_{h}^{*}\nabla\mu_{h}^{n+1}-\mathbf{M}_{h}^{*}\tfrac{\mu_{h}^{*}}{\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1},\tfrac{\omega_{h}}{\vartheta_{h}^{n+1}\vartheta_{h}^{n+1}}\nabla\mu_{h}^{n+1}\rangle$
$\quad+\gamma\langle d^{n+1}_{\tau}\phi_{h}\nabla\phi_{h}^{*},\nabla\omega_{h}\rangle$ (12)
holds for $(\psi_{h},\xi_{h},\omega_{h})\in\mathcal{V}_{h}\times\mathcal{V}_{h}\times\mathcal{V}_{h}$ and all $10\leq n\leq n_{T}-$ 1.
5 Structure-preserving property
In this section we will show that at least one solution to 5 exists and that the numerical method (10)-(12) preserves the structural properties similar to Theorem 4 at the discrete level.
Theorem 6.
Every discrete solution $(\phi_{h},\mu_{h},\vartheta_{h})$ of 5 with $\vartheta_{h}>0$ preserves the structural properties, i.e. the conservation of mass, the discrete energy balance and the entropy production
$\langle d^{n+1}_{\tau}\phi_{h},1\rangle$ $=0,\quad\langle d^{n+1}_{\tau}e_{h},1\rangle=\mathcal{D}_{num}^{n+1}\leq 0,\quad\langle d^{n+1}_{\tau}s_{h},1\rangle=\mathcal{D}_{h}^{n+1}\geq 0.$
The discrete entropy production reads
$\mathcal{D}_{h}^{n+1}=\begin{pmatrix}\tfrac{\nabla\vartheta_{h}^{n+1}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}&\tfrac{\mu_{h}^{*}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1}-\tfrac{\nabla\mu_{h}^{n+1}}{\vartheta_{h}^{n+1}}\end{pmatrix}\begin{pmatrix}\mathbf{K}_{h}^{*}&-\mathbf{C}_{h}^{*}\\ -\mathbf{C}_{h}^{*}&\mathbf{M}_{h}^{*}\end{pmatrix}\begin{pmatrix}\tfrac{\nabla\vartheta_{h}^{n+1}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}\\ \tfrac{\mu_{h}^{*}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1}-\tfrac{\nabla\mu_{h}^{n+1}}{\vartheta_{h}^{n+1}}\end{pmatrix}$
and the numerical dissipation is given as follows
$\mathcal{D}_{num}^{k}=-\tfrac{\tau}{2}$ $\Big{(}\gamma\langle|\nabla d_{\tau}^{k}\phi_{h}|^{2},\vartheta_{h}^{k}\rangle-\langle\partial_{\vartheta\vartheta}f(\phi_{h}^{k-1},\zeta_{1}),(d_{\tau}^{k}\vartheta_{h})^{2}\rangle$
$+\langle\partial_{\phi\phi}f_{\text{vex}}(\zeta_{2},\vartheta_{h}^{k})-\partial_{\phi\phi}f_{\text{cav}}(\zeta_{3},\vartheta_{h}^{k}),(d_{\tau}^{k}\phi_{h})^{2}\rangle\Big{)}.$
Proof.
To derive the conservation of mass and the entropy production, one can proceed exactly as in the continuous case by testing (10) with $\psi_{h}=1$ and (12) with $\omega_{h}=1$, respectively. To get the energy conservation we first use the identity $e=f-\vartheta\cdot\partial_{\vartheta}f$ and calculate as follows
$\langle d^{n+1}_{\tau}e_{h},1\rangle$ $=\langle d^{n+1}_{\tau}f_{h}-d^{n+1}_{\tau}(\vartheta_{h}\cdot\partial_{\vartheta}f_{h}),1\rangle$
$=\tfrac{1}{\tau}\left(\langle f_{h}^{n+1}-f_{h}^{n},1\rangle-\langle\vartheta_{h}^{n+1}\cdot\partial_{\vartheta}f_{h}^{n+1}-\vartheta_{h}^{n}\cdot\partial_{\vartheta}f_{h}^{n},1\rangle\right)$
$=\tfrac{1}{\tau}\left(\langle f_{h}^{n+1}-f_{h}^{n},1\rangle-\langle\partial_{\vartheta}f_{h}^{n+1}-\partial_{\vartheta}f_{h}^{n},\vartheta_{h}^{n+1}\rangle\right.$
$\quad\left.-\langle\partial_{\vartheta}f_{h}^{n},\vartheta_{h}^{n+1}-\vartheta_{h}^{n}\rangle\right)$
$=\tfrac{1}{\tau}\left(\langle f_{h}^{n+1}-f_{h}^{n},1\rangle+\langle s_{h}^{n+1}-s_{h}^{n},\vartheta_{h}^{n+1}\rangle\right.$
$\quad\left.+\tfrac{\gamma}{2}\langle|\nabla\phi_{h}^{n+1}|^{2}-|\nabla\phi_{h}^{n}|^{2},\vartheta_{h}^{n+1}\rangle-\langle\partial_{\vartheta}f_{h}^{n},\vartheta_{h}^{n+1}-\vartheta_{h}^{n}\rangle\right)$
$=\tfrac{1}{\tau}\langle f_{h}^{n+1}-f_{h}^{n},1\rangle-\langle\partial_{\vartheta}f_{h}^{n},d^{n+1}_{\tau}\vartheta_{h}\rangle+\langle d^{n+1}_{\tau}s_{h},\vartheta_{h}^{n+1}\rangle$
$\quad+\gamma\langle\nabla\phi_{h}^{n+1},\vartheta_{h}^{n+1}\nabla d^{n+1}_{\tau}\phi_{h}\rangle-\tfrac{\gamma\tau}{2}\langle|\nabla d^{n+1}_{\tau}\phi_{h}|^{2},\vartheta_{h}^{n+1}\rangle.$
Now we use (12) with $\omega_{h}=\vartheta_{h}^{n+1}$, where the first term cancels out to get
$\langle d^{n+1}_{\tau}e_{h},1\rangle$ $=\tfrac{1}{\tau}\langle f_{h}^{n+1}-f_{h}^{n},1\rangle-\langle\partial_{\vartheta}f_{h}^{n},d^{n+1}_{\tau}\vartheta_{h}\rangle+\gamma\langle\nabla\phi_{h}^{*},d^{n+1}_{\tau}\phi_{h}\nabla\vartheta_{h}^{n+1}\rangle$
$\quad+\langle\tfrac{\mathbf{C}_{h}^{*}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1}+\tfrac{\mathbf{M}_{h}^{*}}{\vartheta_{h}^{n+1}}\nabla\mu_{h}^{n+1}-\mathbf{M}_{h}^{*}\tfrac{\mu_{h}^{*}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1},\nabla\mu_{h}^{n+1}\rangle$
$\quad+\gamma\langle\nabla\phi_{h}^{n+1},\vartheta_{h}^{n+1}\nabla d^{n+1}_{\tau}\phi_{h}\rangle-\tfrac{\gamma\tau}{2}\langle|\nabla d^{n+1}_{\tau}\phi_{h}|^{2},\vartheta_{h}^{n+1}\rangle.$
Insertion of $\xi_{h}=d^{n+1}_{\tau}\phi_{h}$ into (11) shows
$\langle d^{n+1}_{\tau}e_{h},1\rangle$ $=\tfrac{1}{\tau}\langle f_{h}^{n+1}-f_{h}^{n},1\rangle-\langle\partial_{\vartheta}f_{h}^{n},d^{n+1}_{\tau}\vartheta_{h}\rangle+\langle\mu_{h}^{n+1},d^{n+1}_{\tau}\phi_{h}\rangle$
$\quad+\langle\tfrac{\mathbf{C}_{h}^{*}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1}+\tfrac{\mathbf{M}_{h}^{*}}{\vartheta_{h}^{n+1}}\nabla\mu_{h}^{n+1}-\mathbf{M}_{h}^{*}\tfrac{\mu_{h}^{*}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1},\nabla\mu_{h}^{n+1}\rangle$
$\quad-\langle f_{\phi}(\phi_{h}^{n+1},\phi_{h}^{n},\vartheta_{h}^{n+1}),d^{n+1}_{\tau}\phi_{h}\rangle-\tfrac{\gamma\tau}{2}\langle|\nabla d^{n+1}_{\tau}\phi_{h}|^{2},\vartheta_{h}^{n+1}\rangle.$
Finally using (10) with $\psi_{h}=\mu_{h}^{n+1}$ yields
$\langle d^{n+1}_{\tau}e_{h},1\rangle$ $=\tfrac{1}{\tau}\langle f_{h}^{n+1}-f_{h}^{n},1\rangle-\langle\partial_{\vartheta}f_{h}^{n},d^{n+1}_{\tau}\vartheta_{h}\rangle$
$\quad-\langle f_{\phi}(\phi_{h}^{n+1},\phi_{h}^{n},\vartheta_{h}^{n+1}),d^{n+1}_{\tau}\phi_{h}\rangle-\tfrac{\gamma\tau}{2}\langle|\nabla d^{n+1}_{\tau}\phi_{h}|^{2},\vartheta_{h}^{n+1}\rangle.$
Adding ±$\pm f(\phi_{h}^{n},\vartheta_{h}^{n+1})$ and using (A4) allows us the rewrite as follows
$\langle d^{n+1}_{\tau}e_{h},1\rangle$ $=\tfrac{1}{\tau}\langle f_{\text{vex}}(\phi_{h}^{n+1},\vartheta_{h}^{n+1})-f_{\text{vex}}(\phi_{h}^{n},\vartheta_{h}^{n+1}),1\rangle$
$\quad-\langle\partial_{\phi}f_{\text{vex}}(\phi_{h}^{n+1},\vartheta_{h}^{n+1}),d^{n+1}_{\tau}\phi_{h}\rangle$
$\quad+\tfrac{1}{\tau}\langle f_{\text{cav}}(\phi_{h}^{n+1},\vartheta_{h}^{n+1})-f_{\text{cav}}(\phi_{h}^{n},\vartheta_{h}^{n+1}),1\rangle$
$\quad-\langle\partial_{\phi}f_{\text{cav}}(\phi_{h}^{n},\vartheta_{h}^{n+1}),d^{n+1}_{\tau}\phi_{h}\rangle$
$\quad+\tfrac{1}{\tau}\langle f(\phi_{h}^{n},\vartheta_{h}^{n+1})-f(\phi_{h}^{n},\vartheta_{h}^{n}),1\rangle-\langle\partial_{\vartheta}f(\phi_{h}^{n},\vartheta_{h}^{n}),d^{n+1}_{\tau}\vartheta_{h}\rangle$
$\quad-\tfrac{\gamma\tau}{2}\langle|\nabla d^{n+1}_{\tau}\phi_{h}|^{2},\vartheta_{h}^{n+1}\rangle.$
Taylor-expansion of $f$ in $\phi$ and $\vartheta$ finally reveals
$\langle d^{n+1}_{\tau}e_{h},1\rangle$ $=-\tfrac{\tau}{2}\Big{(}\langle\partial_{\phi\phi}f_{\text{vex}}(\zeta_{2},\vartheta_{h}^{n+1})-\partial_{\phi\phi}f_{\text{cav}}(\zeta_{3},\vartheta_{h}^{n+1}),(d_{\tau}^{n+1}\phi_{h})^{2}\rangle$
$\quad-\langle\partial_{\vartheta\vartheta}f(\phi_{h}^{n},\zeta_{1}),(d_{\tau}^{n+1}\vartheta_{h})^{2}\rangle+\gamma\langle|\nabla d_{\tau}^{n+1}\phi_{h}|^{2},\vartheta_{h}^{n+1}\rangle\Big{)}=\mathcal{D}_{num}^{n+1}.$
The negativity of $\mathcal{D}_{num}^{k}$ follows from the properties of $f$ in (A4) and the positivity of $\vartheta_{h}$. ∎
5.1 Existence of positive solutions
In the next step we will establish existence of discrete solutions with uniformly positive temperature. To this end we need to enhance our assumptions on the driving potential $f$.
  1. (A5)
[…]
    $\langle e(\phi,\vartheta)-c_{f}s(\phi,\vartheta),1\rangle\leq C_{1}\Longrightarrow\|\nabla\phi\|_{L^{2}}^{2}+\|\vartheta\|_{L^{1}(\Omega)}\leq C_{*}C_{1}.$
  2. (A6)
    We assume that $|f_{\phi}(\phi_{h}^{n+1},\phi_{h}^{n},\vartheta_{h}^{n+1})|\leq C(|\phi_{h}^{n+1}|^{p_{1}}+|\phi_{h}^{n+1}|^{p_{2}}+|\vartheta_{h}^{n+1}|^{k}+1)$, with $p_{1},p_{2}\leq 6$ and $k\in\mathbb{N}$.
  3. (A7)
    The mesh $\mathcal{T}_{h}$ is quasi-uniform.
Having quasi-uniform meshes $\mathcal{T}_{h}$ we can apply the inverse inequalities, see [6, Theorem 4.5.11],
$\|v_{h}\|_{H^{1}}$ $\leq c_{\text{inv}}h^{-1}\|v_{h}\|_{L^{2}},$ (13)
$\|v_{h}\|_{L^{p}}$ $\leq c_{\text{inv}}h^{d/p-d/q}\|v_{h}\|_{L^{q}},$ (14)
which hold for all discrete functions $v_{h}\in\mathcal{V}_{h}$ and $1\leq q\leq p\leq\infty$.
Theorem 7.
[…]
Proof.
To prove the existence result it suffices to consider a single time step. For simplicity we consider $g^{*}=g^{n}$ for all $g\in\{\phi_{h},\mu_{h}\vartheta_{h}\}$ in 5, but the whole proof works for other choices of $g^{*}$. After choosing the Lagrange basis in $\mathcal{V}_{h}$ we can rewrite 5 as a finite-dimensional nonlinear problem $\mathbf{g}(\mathbf{x})=\mathbf{0}$, where $\mathbf{x}=(\boldsymbol{\phi},\boldsymbol{\mu},\boldsymbol{\vartheta})\in\mathbb{R}^{3N}$ is a vector containing the nodal values of $x_{h}:=(\phi^{n+1}_{h},\mu_{h}^{n+1},\vartheta_{h}^{n+1})$. For simplicity we will use the abbreviation $(\phi_{h},\mu_{h},\vartheta_{h})\equiv(\phi_{h}^{n+1},\mu_{h}^{n+1},\vartheta_{h}^{n+1})$. To apply Theorem 1 we consider
$V$ :=$:=\{(\boldsymbol{\phi},\boldsymbol{\mu},\boldsymbol{\vartheta})\in\mathbb{R}^{3N}\text{ such that }\boldsymbol{\vartheta}>0\},$ (15)
$W$ :=$:=\{(\boldsymbol{\phi},\boldsymbol{\mu},\boldsymbol{\vartheta})\in\mathbb{R}^{3N}\text{ such that }\|\boldsymbol{\phi}\|+\|\boldsymbol{\mu}\|\leq C_{\phi,\mu},\varepsilon<\boldsymbol{\vartheta}<C_{\vartheta},\}.$ (16)
[…]
Extended problem and condition (i):
In the next step we will define $\mathbf{G}(\mathbf{x},\alpha)=\langle T_{\alpha}x_{h},y_{h}\rangle$ with $\langle T_{\alpha}x_{h},y_{h}\rangle$ given by the variational problem
$\frac{1}{\tau}\langle\phi_{h}-\phi_{h}^{n},\psi_{h}\rangle-\alpha\langle\tfrac{\mathbf{M}_{h}^{n}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{n}}\mu_{h}^{n}\nabla\vartheta_{h}-\tfrac{\mathbf{M}_{h}^{n}}{\vartheta_{h}}\nabla\mu_{h}-\tfrac{\mathbf{C}_{h}^{n}}{\vartheta_{h}\cdot\vartheta_{h}^{n}}\nabla\vartheta_{h},\nabla\psi_{h}\rangle=0$ (17)
$\langle\mu_{h},\xi_{h}\rangle-\gamma\langle\nabla\phi_{h},\vartheta_{h}\nabla\xi_{h}\rangle-\gamma\langle\nabla\phi_{h}^{n},\xi_{h}\nabla\vartheta_{h}\rangle-\langle\partial_{\phi}f_{\phi}(\phi_{h},\phi_{h}^{n},\vartheta_{h}),\xi_{h}\rangle=0$ (18)
$\frac{1}{\tau}\langle s(\phi_{h},\vartheta_{h})-s(\phi_{h}^{n},\vartheta_{h}^{n}),\omega_{h}\rangle-\alpha\langle\tfrac{\mathbf{K}_{h}^{n}}{\vartheta_{h}\cdot\vartheta_{h}^{n}}\nabla\vartheta_{h}+\tfrac{\mathbf{C}_{h}^{n}}{\vartheta_{h}}\nabla\mu_{h}-2\tfrac{\mathbf{C}_{h}^{n}}{\vartheta_{h}\cdot\vartheta_{h}^{n}}\mu_{h}^{n}\nabla\vartheta_{h}$
$\quad-\mathbf{M}_{h}^{n}\tfrac{\mu_{h}^{n}}{\vartheta_{h}}\nabla\mu_{h}+\mathbf{M}_{h}^{n}\tfrac{\left(\mu_{h}^{n}\right)^{2}}{\vartheta_{h}\cdot\vartheta_{h}^{n}}\nabla\vartheta_{h},\tfrac{\omega_{h}}{\vartheta_{h}\cdot\vartheta_{h}^{n}}\nabla\vartheta_{h}-\tfrac{1}{\vartheta_{h}^{n}}\nabla\omega_{h}\rangle$
$\quad-\alpha\langle\tfrac{\mathbf{C}_{h}^{n}}{\vartheta_{h}\cdot\vartheta_{h}^{n}}\nabla\vartheta_{h}+\tfrac{\mathbf{M}_{h}^{n}}{\vartheta_{h}}\nabla\mu_{h}-\tfrac{\mathbf{M}_{h}^{n}}{\vartheta_{h}\cdot\vartheta_{h}^{n}}\mu_{h}^{n}\nabla\vartheta_{h},\tfrac{\omega_{h}}{\vartheta_{h}}\nabla\mu_{h}\rangle$
$\quad+\frac{1}{\tau}\gamma\langle\nabla\phi_{h}^{n}(\phi_{h}-\phi_{h}^{n}),\nabla\omega_{h}\rangle=0$ (19)
for all $y_{h}=(\psi_{h},\xi_{h},\omega_{h})\in\mathcal{V}_{h}\times\mathcal{V}_{h}\times\mathcal{V}_{h}$. By construction it holds $\mathbf{G}(\cdot,1)=\mathbf{g}(\cdot)=0$, hence condition $(i)$ of Theorem 1 is fulfilled.
Continuity of the map and condition (ii)
In this part we show that for every $\alpha\in[0,1]$, if a solution $\mathbf{x}\in V$ of $\mathbf{G}(\mathbf{x},\alpha)=0$ exists, then $\mathbf{x}\in W.$ To provide the required bounds of $\|\mathbf{x}\|,\leavevmode\nobreak\ \mathbf{x}=(\boldsymbol{\phi},\boldsymbol{\mu},\boldsymbol{\vartheta})$, we follow the lines of the proof of Theorem 6 and find that for every $\alpha\in[0,1]$ we have
$\langle e(\phi_{h},\vartheta_{h})-e(\phi_{h}^{n},\vartheta_{h}^{n}),1\rangle$ $=\tau\mathcal{D}_{num}^{n+1}\leq 0,$
$\langle s(\phi_{h},\vartheta_{h})-s(\phi_{h}^{n},\vartheta_{h}^{n}),1\rangle$ $=\tau\alpha\mathcal{D}_{h}^{n+1}\geq 0.$
Combination yields
$\langle e(\phi_{h},\vartheta_{h})-c_{f}s(\phi_{h},\vartheta_{h}),1\rangle$ $-\tau\mathcal{D}_{num}^{n+1}+c_{f}\tau\alpha\mathcal{D}_{h}^{n+1}$
$=\langle e(\phi_{h}^{n},\vartheta_{h}^{n})-c_{f}s(\phi_{h}^{n},\vartheta_{h}^{n}),1\rangle\leq C^{\prime}.$
First, since $(\phi_{h},\mu_{h},\vartheta_{h})\in V$ we find that $\vartheta_{h}>0$ and thus there exists $c_{1},c_{2}$ such that $c_{1}\leq\vartheta_{h}\leq c_{2}.$ Note that the lower and upper bounds may depend on $\alpha$. To obtain independent bounds we take the minimum/maximum with respect to $\alpha.$ Furthermore using $\psi_{h}=1$ we derive conservation of mass and hence control on the mean value of $\phi_{h}.$ Using (A5) and Poincare inequality for $\phi_{h}$ we obtain
$\|\phi_{h}\|_{H^{1}}^{2}+\|\vartheta_{h}\|_{L^{1}(\Omega)}\leq C^{*}C^{\prime},$ (20)
where $C^{\prime}$ only depends on the norms of known functions $(\phi_{h}^{n},\vartheta_{h}^{n}).$ It remains to obtain an estimate on $\mu_{h}$:
$\|\mu_{h}\|_{L^{2}}^{2}$ $=\gamma\langle\vartheta_{h}\nabla\phi_{h},\nabla\mu_{h}\rangle+\gamma\langle\nabla\phi_{h}^{n}\nabla\vartheta_{h},\mu_{h}\rangle+\langle f_{\phi}(\phi_{h},\phi_{h}^{n},\vartheta_{h}),\mu_{h}\rangle$
$\leq\gamma c_{2}\|\nabla\phi_{h}\|_{L^{2}}\|\nabla\mu_{h}\|_{L^{2}}+\gamma\|\nabla\phi_{h}^{n}\|_{L^{2}}\|\nabla\vartheta_{h}\|_{L^{2}}\|\mu_{h}\|_{L^{\infty}}$
$\;+\|f_{\phi}(\phi_{h}^{n+1},\phi_{h}^{n},\vartheta_{h}^{n+1})\|_{L^{1}}\|\mu_{h}\|_{L^{\infty}}$
$=(a)+(b)+(c).$
Using the inverse inequality (13) we find that there exists a constant $C=C(h^{-1})$ such that
$(a)$ $\leq C(h^{-1})\|\nabla\phi_{h}\|_{L^{2}}\|\mu_{h}\|_{L^{2}},$
$(b)$ $\leq C(h^{-1})\|\nabla\phi_{h}^{n}\|_{L^{2}}\|\vartheta_{h}\|_{L^{1}}\|\mu_{h}\|_{L^{2}},$
$(c)$ $\leq C(h^{-1})\|f_{\phi}(\phi_{h}^{n+1},\phi_{h}^{n},\vartheta_{h}^{n+1})\|_{L^{1}}\|\mu_{h}\|_{L^{2}}.$
For the last term we use the lower and upper bounds for $\vartheta_{h}$ and (A6) to estimate
$(c)\leq C(\|\phi_{h}\|_{H^{1}}^{p_{1}}+\|\phi_{h}^{n}\|_{H^{1}}^{p_{2}}+\|\vartheta_{h}\|_{L^{1}}^{k}+1)\|\mu_{h}\|_{L^{2}}.$
Using (a), (b), (c), the lower/upper bounds for $\vartheta$ and the bounds (20) we find
$\|\mu_{h}\|_{L^{2}}\leq\bar{C}.$ (21)
Hence, we choose the constants $C_{\phi,\mu},\varepsilon,C_{\vartheta}$ in (16) as
[…]
Condition (iii): Topological degree of $\mathbf{G}(\cdot,0)$:
In this part we show that the problem $\mathbf{G}(\mathbf{x},0)=0$ has a solution $\mathbf{x}\in W$ and that the topological degree of $\mathbf{G}(\cdot,0)$ is non-zero. We start by reformulating $\mathbf{G}(\mathbf{x},0)$, i.e. (17)–(19) with $\alpha=0$, which reads
$\langle\phi_{h}-\phi_{h}^{n},\psi_{h}\rangle$ $=0$ (22)
$\langle\mu_{h},\xi_{h}\rangle-\gamma\langle\nabla\phi_{h},\vartheta_{h}\nabla\xi_{h}\rangle-\gamma\langle\nabla\phi_{h}^{n},\xi_{h}\nabla\vartheta_{h}\rangle-\langle\partial_{\phi}f_{h},\xi_{h}\rangle$ $=0$ (23)
$\langle s(\phi_{h},\vartheta_{h})-s(\phi_{h}^{n},\vartheta_{h}^{n}),\omega_{h}\rangle$ $=0.$ (24)
[…]
Next, we will show that the topological degree is non-zero. To this end we compute the Jacobian of $\mathbf{G}(\cdot,0)$
$\mathbf{J}(\mathbf{G}(\cdot,0))=\begin{pmatrix}\mathbf{A}&\mathbf{0}&\mathbf{0}\\ \mathbf{H}^{\phi}&\mathbf{A}&\mathbf{H}^{\theta}\\ \mathbf{S}^{\phi}&\mathbf{0}&\mathbf{S}^{\vartheta}\end{pmatrix}$
with the matrices
$\mathbf{A}_{ij}$ $=\langle\lambda_{j},\lambda_{i}\rangle,\quad\mathbf{H}_{i,j}^{\phi}=\gamma\langle\nabla\lambda_{j},\vartheta_{h}\nabla\lambda_{i}\rangle+\langle\partial_{\phi_{h}}f_{\phi}(\phi_{h},\phi_{h}^{n},\vartheta_{h})\lambda_{j},\lambda_{i}\rangle,$
$\mathbf{H}_{i,j}^{\theta}=$ $\gamma\langle\nabla\phi_{h},\lambda_{j}\nabla\lambda_{i}\rangle+\gamma\langle\nabla\phi_{h}^{n},\lambda_{i}\nabla\lambda_{j}\rangle+\langle\partial_{\vartheta_{h}}f_{\phi}(\phi_{h},\phi_{h}^{n},\vartheta_{h})\lambda_{j},\lambda_{i}\rangle,$
$\mathbf{S}^{\phi}_{i,j}$ $=\langle\partial_{\phi}s(\phi_{h},\vartheta_{h})\lambda_{j},\lambda_{i}\rangle,\quad\mathbf{S}^{\vartheta}_{i,j}=\langle\partial_{\vartheta}s(\phi_{h},\vartheta_{h})\lambda_{j},\lambda_{i}\rangle.$
Here the functions $\lambda_{i}$ are the finite element basis functions. The topological degree involves the signum of the determinant of the Jacobian. We can compute that $\det(\mathbf{J}(\mathbf{G}(\cdot,0)))=\det(\mathbf{A})^{2}\det(\mathbf{S}^{\vartheta}).$ The mass matrix $\mathbf{A}$ is positive definite, hence $\det(\mathbf{A})^{2}>0$. Since $\partial_{\vartheta}s=-\partial_{\vartheta\vartheta}F$, we find using strict concavity assumptions of $f$, i.e (A5), that $-\partial_{\vartheta\vartheta}F>0$ and therefore $\mathrm{sgn}(\det(\mathbf{J}(\mathbf{G}(\mathbf{x},0))))=1$. The topological degree is given by
$\mathrm{deg}(\mathbf{G}(\cdot,0)):=\sum_{\mathbf{x}\in\mathbf{G}^{-1}(0)}\mathrm{sgn}(\det(\mathbf{J}(\mathbf{G}(\mathbf{x},0))))=\sum_{\mathbf{x}\in\mathbf{G}^{-1}(0)}1.$
Hence, the topological degree is non-zero, if at least one $\mathbf{x}\in\mathbf{G}^{-1}(0)$ exists. We have already shown that system (22)–(24) admits at least one solution. Hence, the topological degree is non-zero.
Finally, we can apply Theorem 1 for one time step of the problem $\mathbf{G}(\mathbf{x},\alpha)=0$ and find that at least one solution $\mathbf{x}\in W$ exists. The existence of at least one discrete solution $x_{h}=(\phi_{h},\mu_{h},\vartheta_{h})$ can be done in an inductive way. ∎
Remark 8.
We should note that Theorem 7 gives the existence of a bounded solution with strictly positive temperature. The bounds depend on $h,\tau$ due to the use of inverse inequalities. Note that the quasi-uniformity assumptions is only for simplicity and can be omitted, if local mesh sizes are considered.
6 Numerical experiments
In this section, we present numerical results for the NCH system (1), (2), (2.3), using our numerical scheme (10)-(12). We consider a convergence test as well as an example that illustrates the schemes behavior. We choose a domain $\Omega=[0,1]^{2}$ with periodic boundary conditions. We consider an example of a quenching process, where the solution starts in a mixed state and is then quenched below the critical temperature $\vartheta_{c}$ at the corners by setting a low initial temperature. At the rest of the domain $\Omega$, the temperature is above $\vartheta_{c}$. To model the temperature dependent separation of two phases, we introduce the driving potential
$f$ $=a\big{(}2\phi^{4}-4\phi^{3}+(\tfrac{\vartheta-\vartheta_{c}}{d}+3)\phi^{2}-(\tfrac{\vartheta-\vartheta_{c}}{d}+1)\phi+0.25(\tfrac{\vartheta-\vartheta_{c}}{d}+1)\big{)}$
$\quad-b(\vartheta\log(\tfrac{\vartheta}{\vartheta_{c}})+\vartheta-\vartheta_{c})$ (25)
with parameters $a=0.01,\leavevmode\nobreak\ b=1,\leavevmode\nobreak\ d=1$ and $\vartheta_{c}=3$. The graphs are depicted in Figure 1. Mobility and heat conductivity are $\mathbf{M}=10^{-2}\cdot\mathbf{I}$ and $\mathbf{K}=5\cdot 10^{-2}\cdot\mathbf{I}$. The cross-coupling term $\mathbf{C}$ will be either $10^{-4}\cdot\mathbf{I}$ or $0\cdot\mathbf{I}$. The interface parameter is fixed to $\gamma=10^{-3}$. For simplicity we take $g^{*}$ in (10)–(12) are taken to be $g^{n}$. It means that they are evaluated explicitly and we set $\mu_{h}^{0}=0$. The resulting system is still nonlinear and solved using the Newton method with error tolerance $10^{-12}$. The resulting linear system is solved using a direct solver. The numerical method is implemented in NGSolve [36].
$\phi$$f(\phi,\vartheta)$$f(\phi,4)$
$\phi$$f(\phi,\vartheta)$$f(\phi,2)$
Figure 1: The driving potential $f$ for different $\vartheta$
6.1 Convergence test
First we want to present experimental convergence results and start with the convergence in space. To this end, we consider the initial data
$\phi^{0}(x,y)$ $:=0.6,$
$\vartheta^{0}(x,y)$ $:=-5.9\cdot\operatorname{trans}_{[0,1]}\left(\frac{\sqrt{0.3x^{2}+0.3y^{2}}-0.2)}{\sqrt{0.001}}\right)+6.0,$
where
$\operatorname{trans}_{[0,1]}(z)=\frac{\min\Big{(}\max\big{(}\tanh(z),-1\big{)},1\Big{)}+1}{2}\in[0,1].$
Since no exact solution is available, we will study self-convergence by computing the errors of the Cauchy sequence. Specifically, the error quantities we consider in space are given as
$\mathrm{err}(\nabla\phi,h_{k})$ $:=\max_{n\in\mathcal{I}_{\tau}}\|\phi^{n}_{h_{k}}-\phi^{n}_{h_{k+1}}\|_{1}^{2},\quad\mathrm{err}(\nabla\mu,h_{k}):=\tau\sum_{n=1}^{n_{\tau}}\|\mu^{n}_{h_{k}}-\mu^{n}_{h_{k+1}}\|_{1}^{2},$
$\mathrm{err}(\vartheta,h_{k})$ $:=\max_{n\in\mathcal{I}_{\tau}}\|\vartheta^{n}_{h_{k}}-\vartheta^{n}_{h_{k+1}}\|_{0}^{2},\quad\mathrm{err}(\nabla\vartheta,h_{k}):=\tau\sum_{n=1}^{n_{\tau}}\|\vartheta^{n}_{h_{k}}-\vartheta^{n}_{h_{k+1}}\|_{1}^{2}.$
To compute the experimental order of convergence (eoc) in space we introduce
$\mathrm{eoc}_{k}:=\log_{2}\left(\frac{\mathrm{err}(a,h_{k-1})}{\mathrm{err}(a,h_{k})}\right)$
for $a\in\{\nabla\phi,\nabla\mu,\vartheta,\nabla\vartheta\}.$
To this end, we consider mesh refinements with the following mesh sizes $h_{k}\approx 2^{-k}$ for $k=4,\ldots,8$ and for the time step we choose $\tau=2^{-10}\cdot 10^{-1}$. The final time is set to $T=10^{-1}$. The errors and approximate convergence rates, with $\mathbf{C}=10^{-4}\cdot\mathbf{I}$, can be found in Table 1. They indicate optimal first-order convergence in the $H^{1}$-norm and the second-order convergence in the $L^{2}$-norm.
Table 1: Errors and rates for convergence in space.
$k$ $\mathrm{err}(\nabla\phi,h_{k})$ $\mathrm{eoc}_{k}$ $\mathrm{err}(\nabla\mu,h_{k})$ $\mathrm{eoc}_{k}$ $\mathrm{err}(\vartheta,h_{k})$ $\mathrm{eoc}_{k}$ $\mathrm{err}(\nabla\vartheta,h_{k})$ $\mathrm{eoc}_{k}$
4 $1.36\cdot 10^{-3}$ $8.37\cdot 10^{-5}$ $5.62\cdot 10^{-3}$ $3.88\cdot 10^{0}$
5 $4.35\cdot 10^{-4}$ $1.65$ $3.17\cdot 10^{-5}$ $1.40$ $2.96\cdot 10^{-4}$ $4.24$ $9.85\cdot 10^{-1}$ $1.98$
6 $1.25\cdot 10^{-4}$ $1.79$ $7.26\cdot 10^{-6}$ $2.13$ $1.90\cdot 10^{-5}$ $3.96$ $2.54\cdot 10^{-1}$ $1.96$
7 $3.21\cdot 10^{-5}$ $1.97$ $1.62\cdot 10^{-6}$ $2.16$ $1.27\cdot 10^{-6}$ $3.91$ $6.25\cdot 10^{-2}$ $2.02$
For the time convergence, the expected optimal order is first-order convergence. Therefore, we use time step refinements $\tau_{k}=2^{-k}\cdot 10^{-1}$ for $k=7,...,11$ and simulate again until a final time $T=10^{-1}$. The spatial mesh is fixed with mesh size $h=2^{-7}$. Here we consider similar errors as above, namely:
$\text{err}(\nabla\phi,\tau_{k})$ $:=\max_{n\in\mathcal{I}_{\tau_{k}}}\|\phi^{n}_{h,\tau_{k}}-\phi^{n}_{h,\tau_{k+1}}\|_{1}^{2},\leavevmode\nobreak\ \text{err}(\nabla\mu,\tau_{k}):=\tau_{k}\sum_{n=1}^{n_{\tau_{k}}}\|\mu^{n}_{h,\tau_{k}}-\mu^{n}_{h,\tau_{k+1}}\|_{1}^{2},$
$\text{err}(\vartheta,\tau_{k})$ $:=\max_{n\in\mathcal{I}_{\tau_{k}}}\|\vartheta^{n}_{h,\tau_{k}}-\vartheta^{n}_{h,\tau_{k+1}}\|_{0}^{2},\leavevmode\nobreak\ \text{err}(\nabla\vartheta,\tau_{k}):=\tau_{k}\sum_{n=1}^{n_{\tau_{k}}}\|\vartheta^{n}_{h,\tau_{k}}-\vartheta^{n}_{h,\tau_{k+1}}\|_{1}^{2}.$
To compute the experimental order of convergence (eoc) in time we introduce
$\mathrm{eoc}_{k}:=\log_{2}\left(\frac{\mathrm{err}(a,\tau_{k-1})}{\mathrm{err}(a,\tau_{k})}\right)$
for $a\in\{\nabla\phi,\nabla\mu,\vartheta,\nabla\vartheta\}$. The errors and rates, for $\mathbf{C}=10^{-4}\cdot\mathbf{I}$, can be found in Table 2. As expected the optimal first-order convergence in time is achieved.
Table 2: Errors and rates for convergence in time.
$k$ $\text{err}(\nabla\phi,\tau_{k})$ $\mathrm{eoc}_{k}$ $\text{err}(\nabla\mu,\tau_{k})$ $\mathrm{eoc}_{k}$ $\text{err}(\vartheta,\tau_{k})$ $\mathrm{eoc}_{k}$ $\text{err}(\nabla\vartheta,\tau_{k})$ $\mathrm{eoc}_{k}$
7 $2.08\cdot 10^{-7}$ $1.72$ $7.97\cdot 10^{-9}$ $1.94$ $4.43\cdot 10^{-8}$ $2.01$ $1.25\cdot 10^{-5}$ $2.01$
8 $5.89\cdot 10^{-8}$ $1.82$ $2.04\cdot 10^{-9}$ $1.97$ $1.11\cdot 10^{-8}$ $2.00$ $3.11\cdot 10^{-6}$ $2.01$
9 $1.59\cdot 10^{-8}$ $1.89$ $5.16\cdot 10^{-10}$ $1.98$ $2.76\cdot 10^{-9}$ $2.00$ $7.76\cdot 10^{-7}$ $2.00$
10 $4.15\cdot 10^{-9}$ $1.94$ $1.30\cdot 10^{-10}$ $1.99$ $6.90\cdot 10^{-10}$ $2.00$ $1.94\cdot 10^{-7}$ $2.00$
Similar convergence results in space and time can be obtained without the cross-coupling term, i.e. $\mathbf{C}=0\cdot\mathbf{I}$.
6.2 Illustrating example
In this subsection we illustrate the behavior of the NCH system. We keep $\theta^{0}$ as before and choose
$\phi^{0}(x,y)$ $:=0.5+0.01\sin(211\pi x)\sin(211\pi y).$
All parameters are choosen as above. We work with the mobility matrix $\mathbf{M}=\mathbf{I}$. Furthermore, we will compare the results obtained for $\mathbf{C}=0\cdot\mathbf{I}$ and $\mathbf{C}=10^{-4}\cdot\mathbf{I}$, i.e. with and without cross-coupling.
The temporal evolution of $\phi_{h}$ for $\mathbf{C}=0\cdot\mathbf{I}$ and $\mathbf{C}=10^{-4}\cdot\mathbf{I}$ is depicted in Figure 2. The results of $\vartheta_{h}$ can be found in Figure 3.
$t=0.1$ $t=0.5$ $t=2$ $t=10$
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Snapshots of volume fraction $\phi$ for different cross-coupling matrices $\mathbf{C}$: (Top) $\mathbf{C}=0\cdot\mathbf{I}$; (Bottom) $\mathbf{C}=10^{-4}\cdot\mathbf{I}$.
$t=0.1$ $t=0.5$ $t=2$ $t=10$
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Snapshots of volume fraction $\vartheta$ for different cross-coupling matrices $\mathbf{C}$: (Top) $\mathbf{C}=0\cdot\mathbf{I}$; (Bottom) $\mathbf{C}=10^{-4}\cdot\mathbf{I}$.
We observe that in both cases the temperature profile is essentially the same. Due to the heat conduction, the sharp temperature gradients of the initial condition are smeared and heat dissipates into the colder regions. The temperature remains above the critical temperature $\vartheta_{c}$ in the middle of the domain and nearly zero at the corners, i.e. below the critical temperature. For the phase variable $\phi_{h}$, we see that for $\mathbf{C}=0\cdot\mathbf{I}$ a local phase separation occurs. This means that in a certain area around the four corners the mixture starts to separate into distinct phases. In the case of $\mathbf{C}=10^{-4}\cdot\mathbf{I}$, where we allow cross-coupling, the initial evolution is comparable, however we can observe more separation at the interface of two phases and also in the middle of the domain. This is done due to the additional driving force in the Cahn-Hilliard equation, leading to a final state with a fully connected one phase.
Refer to caption Refer to caption
Figure 4: Structure preserving properties over time (Top left): Mass conservation error; (Top right): Energy dissipation; (Bottom) Entropy production
Structure-preserving properties, i.e. the conservation of mass, and energy dissipation as well as the entropy production, are shown in Figure 4 for both cases. Numerical results confirm theoretical results from Section 5. Specifically, mass conservation up to a numerical error of $10^{-14}$, energy dissipation of order $10^{-3}$ as well as the entropy production. We note that the entropy production in both cases only differ by a factor of $10^{-4}$.
7 Conclusion & Outlook
In this work we have proposed a fully discrete finite element method for the non-isothermal Cahn-Hilliard equation based on the use of the entropy equation instead of the internal energy equation. The scheme preserves the conservation of mass and the entropy production, while the internal energy, here the same as total energy, is dissipated in time due to numerical diffusion. Numerical experiments document optimal order of convergence, i.e. the first order in time and the first or second order in space depending on the norm. We also illustrate the impact of the cross-coupling terms between the Cahn-Hilliard and the entropy equation by numerical simulations. In future work we want to extend our scheme to eliminate the numerical dissipation such that the discrete energy is conserved. Furthermore, we will consider extensions to incompressible flows.
Acknowledgement
The present research has been support by the German Science Foundation under the grant TRR146 Multiscale Simulation Methods for Soft Matter Systems and by the grant SPP 2256 Variational Methods for Predicting Complex Phenomena in Engineering Structures and Materials. M.L.-M. gratefully acknowledges the support of the Gutenberg Research Fellowship and the Mainz Institute for Multiscale Modelling (M3odel).